Calculate the population variance of the heights of trees(in meters)




10.1, 10.2, 12.4, 3.6,14.3,15.9,16.17

1
Expert's answer
2021-12-09T14:44:08-0500
"mean=\bar{x}=\dfrac{1}{n}\sum_ix_i=\dfrac{1}{7}(10.1+10.2"

"+ 12.4+ 3.6+14.3+15.9+16.17)"

"=\dfrac{82.67}{7}=11.81"

"Var(X)=\sigma^2=\dfrac{1}{n}\sum_i(x_i-\bar{x})^2"

"=\dfrac{1}{7}((10.1-11.81)^2+(10.2-11.81)^2"

"+(12.4-11.81)^2+(3.6-11.81)^2"

"+(14.3-11.81)^2+(15.9-11.81)^2"

"+(16.17-11.81)^2)=\dfrac{115.2062}{7}\approx16.4580"